# Mitigating Covariate Shift in Imitation Learning for Autonomous Vehicles Using Latent Space Generative World Models

### v3

@@ III-C Training and Losses
We use variational inference approach to train the system by maximizing the ELBO [22]. Our objective function is , similar to [12], where
@@ IV-A CARLA Simulator Closed-loop Evaluation
Our trained system was deployed in the closed-loop CARLA simulator [6] for evaluation. Our system successfully reacted to traffic lights, followed lead vehicles, stopped for them, avoided collisions, allowed pedestrians to pass, and adhered to a given route. All training and inference were performed end-to-end without any handcrafted code. We quantitatively evaluated our system by computing closed-loop metrics according to the CARLA Leaderboard 1.0 (Sensors Track).222Like MILE [12], we were unable to submit to the online leaderboard, https://leaderboard.carla.org/leaderboard which is now closed. CARLA leaderboard 2.0 was not possible, since it would require uploading our code; and it contains few methods. Following MILE [12], we evaluated on held-out Town05 and with new weather conditions, which were excluded from training.
@@ IV-B Covariate Shift Experiments
We compared the two approaches in the closed-loop CARLA simulator using Town02. Results shown in Table II demonstrate successful navigation of our system without drifting into undesirable states. In contrast, the policy trained via behavior cloning (BC) achieves worse scores across multiple metrics in the same simulator.333Adding stochastic sampling to BC improves results slightly, but still leads to drifting and collisions. In our experience, training BC on an order of magnitude more data does not resolve covariate shift. The significant difference demonstrates the effectiveness of our latent space generative world model in mitigating covariate shift. Note, however, that our BC implementation still achieves good performance, primarily due to the robust visual features obtained by DINOv2.

### v4

@@ III-C Training and Losses
We use variational inference approach to train the system by maximizing the ELBO [22].222Note that our ELBO-based loss is the same as the free-energy principle objective used in active inference within cognitive science [27]. Our objective function is , similar to [12], where
@@ IV-A CARLA Simulator Closed-loop Evaluation
Our trained system was deployed in the closed-loop CARLA simulator [6] for evaluation. Our system successfully reacted to traffic lights, followed lead vehicles, stopped for them, avoided collisions, allowed pedestrians to pass, and adhered to a given route. All training and inference were performed end-to-end without any handcrafted code. We quantitatively evaluated our system by computing closed-loop metrics according to the CARLA Leaderboard 1.0 (Sensors Track).333Like MILE [12], we were unable to submit to the online leaderboard, https://leaderboard.carla.org/leaderboard which is now closed. CARLA leaderboard 2.0 was not possible, since it would require uploading our code; and it contains few methods. Following MILE [12], we evaluated on held-out Town05 and with new weather conditions, which were excluded from training.
@@ IV-B Covariate Shift Experiments
We compared the two approaches in the closed-loop CARLA simulator using Town02. Results shown in Table II demonstrate successful navigation of our system without drifting into undesirable states. In contrast, the policy trained via behavior cloning (BC) achieves worse scores across multiple metrics in the same simulator.444Adding stochastic sampling to BC improves results slightly, but still leads to drifting and collisions. In our experience, training BC on an order of magnitude more data does not resolve covariate shift. The significant difference demonstrates the effectiveness of our latent space generative world model in mitigating covariate shift. Note, however, that our BC implementation still achieves good performance, primarily due to the robust visual features obtained by DINOv2.